# Optimisation models for the design of multiple self-consumption loops in semi-rural areas

### v2

@@ Abstract
[inst1]organization=IMT Mines Albi, addressline=1 allée des sciences, Campus Jarlard, city=Albi, postcode=81000, country=France
@@ 1 Introduction
Moreover, these sources have emission points that are much more decentralised and scattered across the territory than traditional sources such as nuclear or coal power plants. All these characteristics imply a degree of uncertainty which impacts national distribution networks.
@@
An overarching decision-making system would be necessary to facilitate the establishment of collective self-consumption communities including solar energy companies, other producing and consuming companies and individual dwellings. This system should consider technical, economical, and legal constraints and propose feasible and optimal loops. Such a system will also facilitate discussions on the potential profitability of various communities. This need for an integrated decision support system is notably formulated by Amarenco - a solar energy production company - and motivated our work to design optimal self-consumption loops that should be investigated.
@@
The next section presents related work on individual and collective self-consumption optimisation with a focus on business aspects and optimisation methods. Section 3 details the single-loop model (SLCpct) and multi-loop model (MLCpct) as well as their Dantzig-Wolfe (MLCol) and Benders (SLExt, MLColExt) decompositions. Section 4 details the creation process of realistic instances to apply the models. This section also defines business indicators built out of models’ variables and parameters that can help evaluate proposed community designs. Section 5 presents results of the application of the models on priory defined realistic instances. The performance of the models is discussed, and insights regarding the application of the models in a business context are provided. Section 6 provides some perspectives regarding the developed models and their potential coupling with domain-related knowledge bases, leading to the conclusion and planned future works.
@@ 2 Related work
A literature review has been conducted under the criteria shown in Table 1, which describes two main aspects of previous studies, one being related to the size of the tackled problems and used methodology, and the other related to business issues. In this systematic literature review, the objective was to find articles using optimisation techniques to ease the building of self-consumption operations that include several actors in a given community.
@@
Table 2 sums up the conducted literature review and compares our approach to some of the analysed articles. The presented articles are selected as they encompass one or several dimensions of our approach.
Table 2: Comparison of analysed articles
@@ 2.1 Specific application context
Related lines of work show that electricity consumption optimisation is an active topic of interest. Despite some studies focusing on the specific cases of multi-energy communities, including gas, wind, and/or heat as energy sources (Bio Gassi and Baysal, , 2022; Gullì et al., , 2022; Pinto et al., , 2020; Bokkisam et al., , 2022; Bahret and Eltrop, , 2021), most of identified papers deal with solar energy only, as it represents the majority of current collective self-consumption projects.
@@
Extending the breadth of considered actors also implies an extension of the considered area. In most cases, proposed algorithms and models are applied to an urban case study (Brusco et al., , 2023; Goitia-Zabaleta et al., , 2023; Gullì et al., , 2022; Mustika et al., 2022a, ; Perger et al., , 2021). Beyond revealing that rural or semi-rural self-consumption loop configurations are rarely considered in the conception of optimisation models, it also underlines the fact that most current self-consumption loops are essentially built in urban context thus limiting the range of actors that can be considered. If urban contexts may favour collective self-consumption as the density of consumers and prosumers is higher than in rural contexts, focusing on semi-rural contexts can also have some advantages, as the constraint of geographic contiguity can be revised by regulatory texts and open to new types of producers. Luz and E Silva, (2021) for example showed the benefits of creating an energy community between a rural winery and the nearby city.
@@
In a nutshell, only few previous lines of work treat the problem on both spatial dimension (large number of actors) and temporal dimension (large time horizon and high number of time steps). Above all, the proposed approaches are mainly focused on flow repartition, and very few of them address the design aspect of collective self-consumption loops definition.
@@ 2.2 Modelling and solving approaches
    Model decompositions to scale to larger problems spatially and regarding time horizon.
@@ 3 Optimisation models
From a business point of view, the design problem can be considered at an operational level (day-to-day optimisation of self-consumption in a small neighbourhood). Our objective though is to treat strategic network design problems, as collective self-consumption is planned for several years, on potentially large territories. This changes the size of studied problems, making it necessary to scale them in time and space through the presented decompositions. The combination of the initial models with their decomposition led to the 5 following models, also shown in Figure 1, whose performance is compared in Section 5:
@@ 3.2 Initial single-loop network design and flow circulation problem
3.2.1 Constraints from self-consumption applications
Definition 3.3 (Self-consumption).
Definition 3.4 (Collective self-consumption).
Individual self-consumption constraints
Collective self-consumption constraints
The interest of participating in collective self-consumption is to favour locally produced electricity instead of grid exchanges. Each net producer must then exchange electricity with other loop members when the loop is globally in need of electricity. Similarly, net consumers must favour electricity produced inside the loop instead of electricity from the grid. In Model (3.2.6), these constraints are expressed through Equations (2e) to (2i) and equivalence (2n).
Limiting energy circulation through actors
The total amount of energy distributed by an actor (collective self-consumption + energy sold to the grid) should not exceed the amount of produced energy. In Model (3.2.6), this constraint is expressed with equation (2j). Similarly to distributed energy, the total energy amount that is consumed (collective self-consumption and energy bought to the network) by an actor should not exceed the actor’s consumption, which is expressed as follows:
Actor exclusion constraints
When actor $i$ is excluded from the self-consumption loop, we have $x_{i}=0$. As a consequence of its exclusion, an actor can’t exchange electricity with other actors. It means that $e_{ij}$ variables are subject to the value of $x_{i}$ and $x_{j}$ for each pair of actors, as depicted by Constraints (2b) and (2c) of Model (3.2.6).
3.2.2 Maximal total installed power of a self-consumption loop
To ensure that the loop respects the legal installed power restriction presented previously, Constraint (2m) is set to prevent the global installed power from exceeding the legal criteria.
3.2.3 Neighbourhood graph and distance restriction
We will denote by $\mathcal{E}$ the set of edges in the graph, and $\mathcal{N}_{i}$ the neighbours of node $i$. To cope with the legal restrictions concerning the geographical distance between each pair of actors participating in a self-consumption loop ($D_{ij}<D_{leg}$) equation (2l) refers to $\mathcal{E}$ to avoid non-neighbours to be part of the same collective self-consumption loop and define associated constraints.
3.2.4 Uncertainty in the production and consumption profiles
3.2.5 Objective function
$f_{\mathrm{obj}}=\sum_{i\in[n]}\Bigl{(}F_{i}f_{i}-R_{i}r_{i}\Bigl{)}.$
Fixed inclusion costs ($T_{i}$ adhesion costs of each actor $i$), as well as the price of the electricity within the loop ($r$), are not considered in the objective function, as they will be determined once the geometry of the loop and exchange happening inside the loop are set. Fixing $T$ after this first optimisation step to reach equity between actors involved in the self-consumption loop is evoked as a perspective of this work (Section 6).
3.2.6 Complete single-loop compact model (SLCpct)
All previously defined constraints are aggregated in Model (3.2.6) below. This model expresses the coupled problem of network design and flow repartition in a single-loop configuration, meaning that either belongs to this loop or is not participating in collective self-consumption.
@@ 3.2.6 Complete single-loop compact model (SLCpct)
$e,r,f,y\geq 0$ (2k)
$x_{i}+x_{j}\leq 1$ $\forall(i,j)\notin\mathcal{E}$ (2l)
$\sum_{i\in[n]}x_{i}P^{\text{inst}}_{i}\leq P^{\text{inst}}_{\text{leg}}$ (2m)
$z^{st}=1\Leftrightarrow\sum_{i\in[n]}x_{i}(P_{i}^{st}-C_{i}^{st})\geq 0$ $\forall s\in\mathcal{S},t\in\mathcal{T}$ (2n)
$x\in\left\{0,1\right\}^{n},z\in\left\{0,1\right\}^{|\mathcal{S}|\,|\mathcal{T}|}.$ (2o)
@@ 3.3 Benders decomposition for larger time scales (SLExt)
We can use a Benders decomposition of the problem, expressing the flow optimality conditions as constraints on the $x$, $w$, and $z$ variables. Benders decomposition exploits the separation of variables between the binding variables which design the network and the continuous variables that are fully block-separable (see Rahmaniani et al., (2017) for a recent literature review of the Benders decomposition, applications and algorithms). We can note $\phi_{st}:\{0,1\}^{n}\times\{0,1\}\mapsto\mathbb{R}$ the optimal value function of the problem given fixed values of $x$ and $z^{st}$. Model (3.2.6) is equivalent to:
@@
The feasible region of Model (3.3) being independent of $x$ and $z$, we can replace Constraint (3b) with:
@@ 3.4 Introduction of multiple loops (MLCpct)
Aside of $x_{i}^{l}$ variables, some of the constraints defined in Section 3.2 evolve as soon as multiple loops are considered. The exclusion constraints (Constraints (2b) and (2c)) should be rewritten as follows, to ensure that two actors not being members of the same loop can’t exchange electricity:
@@
where $D_{i,j}=min(P_{i},C_{j})$. To define this constraint (Constraint (7l) in Model (3.4)), we introduced $w_{ij}^{l}$ which are binary variables linearising $x_{i}^{l}x_{j}^{l}$, expressing the membership of both actors $i$ and $j$ to loop $k$ and constraining the flow through (6). These $w_{ij}^{l}$ variables are subject to:
$w_{ij}^{l}\leq x_{i}^{l},\forall(i,j)\in\{1,\ldots n\}^{2},i\neq j,l\in\mathcal{L}$ (Constraint (7m) in Model (3.4))
@@
Legal restriction constraints (Constraints (7h) and  (7i)), collective self-consumption constraints (Constraints (7c) to (7g) and (7p)) and circulation constraint (Constraint (7j)) remain unchanged, and are now indexed by the loop $l$. Additionally, big M constraints are redefined to allow the $y_{i}^{l}$ variables to take any value when actor $i$ is included in loop $l$. These changes result in Model (3.4) which presents symmetries similar to other clustering problems, any permutation of the loop assignments being equivalent. Constraint (7o) reduces the number of symmetric solutions by enforcing solutions with several actors decreasing with the loop index. Constraint (7k) is added to ensure that each actor cannot belong to two different loops.
@@ 3.5 Dantzig-Wolfe decomposition (MLCol)
Assumption 3.6 holds for Model (3.4).
@@
Furthermore, additional business-related constraints can directly be included in the loop, including, for instance, the fact that two coupled actors $(i,j)\in\mathcal{D}$ are both present or both absent from a loop:
@@
The extended formulation of Model (3.4) is essentially captured by the new constraints:
@@ 3.6 Benders decomposition of MLCol model (MLColExt)
s.t. $e_{ij}\leq D_{ij}\sum_{h\in\mathcal{H}}v_{h}a_{i}^{h}a_{j}^{h}$ $(\alpha_{ij})$ (11b)
$\sum_{j\in\mathcal{N}_{i}}e_{ij}+r_{i}-f_{i}-\sum_{j\in\mathcal{N}_{i}}e_{ji}=P_{i}-C_{i}$ $(\pi_{i})$ (11c)
$r_{i}\leq(1-\sum_{h\notin\mathcal{G}}v_{h}a_{i}^{h})Q$ $(\gamma_{i})$ (11d)
$\sum_{i\in[n]}r_{i}a_{i}^{h}\leq\sum_{i\in[n]}a_{i}^{h}v_{h}(P_{i}-C_{i})+Q(1-v_{h})$ $\forall h\in\mathcal{G}\;\;$ $(\nu_{h})$ (11e)
$\sum_{j\in\mathcal{N}_{i}}e_{ij}+r_{i}\leq P_{i}$ $(\tau_{i})$ (11f)
@@ 3.6.1 Dual and cuts
Dualisation of Problem (3.6) gives the following model, whose objective value provides cuts that can be applied to the main model to refine the chosen design of collective self-consumption loops:

### v3

@@ Abstract
[inst1]organization=IRIT, Université de Toulouse, CNRS, Toulouse INP, INU Champollion, addressline=118 Route de Narbonne, city=Toulouse, postcode=31000, state=Occitanie, country=France
@@
[…]
[inst3]organization=IMT Mines Albi, addressline=1 allée des sciences, Campus Jarlard, city=Albi, postcode=81000, country=France
Conflict of Interest
Yohann Chasseray declares that he has no conflict of interest. Mathieu Besançon declares that he has no conflict of interest. Xavier Lorca declares that he has no conflict of interest. Éva Petitdemange declares that she has no conflict of interest.
@@ 1 Introduction
Moreover, these sources have injection points that are much more decentralised and scattered across the territory than traditional sources such as nuclear or coal power plants. All these characteristics imply a degree of uncertainty which impacts national distribution networks.
@@
An overarching decision-making system would be necessary to facilitate the establishment of collective self-consumption communities including solar energy companies, other producing and consuming companies and individual dwellings. This system should consider technical, economical, and legal constraints and propose feasible and optimal loops. Such a system will also facilitate discussions on the potential profitability of various communities. This need for an integrated decision support system is notably formulated by Amarenco - a solar energy production company - and motivated as a real use-case our work to design optimal self-consumption loops.
@@
The next section presents related work on individual and collective self-consumption optimisation with a focus on application-specific modelling choices and optimisation methods. Section 3 details the single-loop model (SLCpct) and multi-loop model (MLCpct) as well as their Dantzig-Wolfe (MLCol) and Benders (SLExt, MLColExt) decompositions. Section 4 details the creation process of realistic instances to apply the models. This section also defines operational indicators built out of model variables and parameters that can help evaluate proposed community designs. Section 5 presents results of the application of the models on priory defined realistic instances. The performance of the models is discussed, and insights regarding the application of the models in the context of the application are provided. Section 6 provides some perspectives regarding the developed models and their potential coupling with domain-related knowledge bases, leading to the conclusion and planned future works.
@@ 2 Related work
A literature review has been conducted under the criteria shown in Table 1, which describes two main aspects of previous studies, one being related to the size of the tackled problems and used methodology, and the other related to application-specific issues. In this systematic literature review, the objective was to find articles using optimisation techniques to ease the building of self-consumption operations that include several actors in a given community.
@@
Table 2 summarizes the conducted literature review and compares our approach to some of the analysed articles. The presented articles are selected as they encompass one or several dimensions of our approach.
Table 2: Comparison of analysed articles (NR: Not Relevant - NA: Not Available)
@@ 2.1 Specific application context
Related lines of work show that electricity consumption optimisation is an active topic of interest. Despite some studies focusing on the specific cases of multi-energy communities, including gas, wind, and/or heat as energy sources (Bio Gassi and Baysal, , 2022; Gullì et al., , 2022; Pinto et al., , 2020; Bokkisam et al., , 2022; Bahret and Eltrop, , 2021), most of the identified papers deal with solar energy only, as it represents the majority of current collective self-consumption projects.
@@
Extending the breadth of considered actors also implies an extension of the considered area. In most cases, proposed algorithms and models are applied to an urban case study (Brusco et al., , 2023; Goitia-Zabaleta et al., , 2023; Gullì et al., , 2022; Mustika et al., 2022a, ; Perger et al., , 2021). Beyond revealing that rural or semi-rural self-consumption loop configurations are rarely considered in the conception of optimisation models, it also underlines the fact that most current self-consumption loops are essentially built in urban context thus limiting the range of actors that can be considered. If urban contexts may favour collective self-consumption as the density of consumers and prosumers is higher than in rural contexts, focusing on semi-rural contexts can also have some advantages, as the constraint of geographic contiguity can be revised by regulatory texts and open to new types of producers. Luz and E Silva, (2021) showed for instance the benefits of creating an energy community between a rural winery and the nearby city.
@@
In a nutshell, only few previous lines of work treat the problem on both a spatial dimension (large number of actors) and a temporal dimension (large time horizon and high number of time steps). Above all, the proposed approaches are mainly focused on flow repartition, and very few of them address the design aspect of collective self-consumption loops definition. Among them, Al-Sorour et al., (2022) proposes models that include the selection of peer-to-peer household pairs, ensuring that each household is involved in only one pairing. Since this approach involves actors selection, the design context differs as the pairing is determined through semi real-time forecasting (selection is conducted on a 2-day ahead forecast), and the model’s complexity lies in its consideration of storage facilities. Moreover, the pairing selection is applied to a predefined and relatively small group of actors (6 households), which minimizes solving time concerns that typically arise in larger territories, where the selection must be performed over a greater number of actors.
@@ 2.2 Modelling and solving approaches
    Model decompositions to scale to larger problems spatially and regarding time horizon. Very few work use both decompositions, as their combination highly complexifies problems. Thus, we provide in this article a simplification of the column generation problem that allows to extend the size of the models both regarding time dimension and spatial dimension.
@@ 3 Optimisation models
From an application point of view, the design problem can be considered at an operational level (day-to-day optimisation of self-consumption in a small neighbourhood). Our objective though is to treat strategic network design problems, as collective self-consumption is planned for several years, on potentially large territories. This changes the size of studied problems, making it necessary to scale them in time and space through the presented decompositions. The combination of the initial models with their decomposition led to the 5 following models, also shown in Figure 1, whose performance is compared in Section 5:
@@ 3.2 Initial single-loop network design and flow circulation problem
The complete single-loop compact model provided as Model (1) (SLCpct) expresses the coupled problem of network design and flow repartition in a single-loop configuration, meaning that each prosumer either belongs to this loop or is not participating in collective self-consumption. The detail about each constraint and the objective function are given throughout section 3.2.1 to 3.2.5.
@@
$x_{i}+x_{j}\leq 1$ $\forall(i,j)\notin\mathcal{E}$ (1k)
$\sum_{i\in[n]}x_{i}P^{\text{inst}}_{i}\leq P^{\text{inst}}_{\text{leg}}$ (1l)
$z^{st}=1\Leftrightarrow\sum_{i\in[n]}x_{i}(P_{i}^{st}-C_{i}^{st})\geq 0$ $\forall s\in\mathcal{S},t\in\mathcal{T}$ (1m)
$x\in\left\{0,1\right\}^{n},z\in\left\{0,1\right\}^{|\mathcal{S}|\,|\mathcal{T}|}$ (1n)
$e,r,f,y\geq 0.$ (1o)
@@
3.2.1 Constraints from self-consumption applications
Definition 3.3 (Self-consumption).
Definition 3.4 (Collective self-consumption).
Individual self-consumption constraints
Collective self-consumption constraints
The interest of participating in collective self-consumption is to favour locally produced electricity instead of grid exchanges. Each net producer must then exchange electricity with other loop members when the loop is globally in need of electricity. Similarly, net consumers must favour electricity produced inside the loop instead of electricity from the grid. In Model (1), these constraints are expressed through Equations (1e) to (1i) and equivalence (1m).
Limiting energy circulation through actors
The total amount of energy distributed by an actor (collective self-consumption + energy sold to the grid) should not exceed the amount of produced energy. In Model (1), this constraint is expressed with equation (1j). Similarly to distributed energy, the total energy amount that is consumed (collective self-consumption and energy bought to the network) by an actor should not exceed the actor’s consumption, which is expressed as follows:
Actor exclusion constraints
When actor $i$ is excluded from the self-consumption loop, we have $x_{i}=0$. As a consequence of its exclusion, an actor cannot exchange electricity with other actors. It means that $e_{ij}$ variables are subject to the value of $x_{i}$ and $x_{j}$ for each pair of actors, as depicted by Constraints (1b) and (1c) of Model (1).
3.2.2 Maximal total installed power of a self-consumption loop
To ensure that the loop respects the legal installed power restriction presented previously, Constraint (1l) is set to prevent the global installed power from exceeding the legal criteria.
3.2.3 Neighbourhood graph and distance restriction
We will denote by $\mathcal{E}$ the set of edges in the graph, and $\mathcal{N}_{i}$ the neighbours of node $i$. To cope with the legal restrictions concerning the geographical distance between each pair of actors participating in a self-consumption loop ($D_{ij}<D_{leg}$) equation (1k) refers to $\mathcal{E}$ to avoid non-neighbours to be part of the same collective self-consumption loop and define associated constraints.
3.2.4 Uncertainty in the production and consumption profiles
3.2.5 Objective function
$f_{\mathrm{obj}}=\sum_{i\in[n]}\Bigl{(}F_{i}f_{i}-R_{i}r_{i}\Bigl{)}.$
Fixed inclusion costs ($T_{i}$ adhesion costs of each actor $i$), as well as the price of the electricity within the loop ($r$), are not considered in the objective function, as they will be determined once the geometry of the loop and exchange happening inside the loop are set. Fixing $T$ after this first optimisation step to reach equity between actors involved in the self-consumption loop is evoked as a perspective of this work (Section 6). Despite balancing the benefits of the actors is left for future work, the proposed model assures a maximal global benefit and a non-zero benefit for each actor involved in the self-consumption loop.
@@ 3.3 Benders decomposition for larger time scales (SLExt)
We can use a Benders decomposition of the problem, expressing the flow optimality conditions as constraints on the $x$, $w$, and $z$ variables. Benders decomposition exploits the separation of variables between the binding variables which design the network and the continuous variables that are fully block-separable (see Rahmaniani et al., (2017) for a recent literature review of the Benders decomposition, applications and algorithms). We can note $\phi_{st}:\{0,1\}^{n}\times\{0,1\}\mapsto\mathbb{R}$ the optimal value function of the problem given fixed values of $x$ and $z^{st}$. Model (1) is equivalent to:
@@
The feasible region of Model (5) being independent of $x$ and $z$, we can replace Constraint (3b) with:
@@ 3.4 Introduction of multiple loops (MLCpct)
Aside of $x_{i}^{l}$ variables, some of the constraints defined in Section 3.2 evolve as soon as multiple loops are considered. The exclusion constraints (Constraints (1b) and (1c)) should be rewritten as follows, to ensure that two actors not being members of the same loop cannot exchange electricity:
@@
where $D_{i,j}=min(P_{i},C_{j})$. To define this constraint (Constraint (7l) in Model (7)), we introduced $w_{ij}^{l}$ which are binary variables linearising $x_{i}^{l}x_{j}^{l}$, expressing the membership of both actors $i$ and $j$ to loop $k$ and constraining the flow through (6). These $w_{ij}^{l}$ variables are subject to:
$w_{ij}^{l}\leq x_{i}^{l},\forall(i,j)\in\{1,\ldots n\}^{2},i\neq j,l\in\mathcal{L}$ (Constraint (7m) in Model (7))
@@
Legal restriction constraints (Constraints (7h) and  (7i)), collective self-consumption constraints (Constraints (7c) to (7g) and (7p)) and circulation constraint (Constraint (7j)) remain unchanged, and are now indexed by the loop $l$. Additionally, big M constraints are redefined to allow the $y_{i}^{l}$ variables to take any value when actor $i$ is included in loop $l$. These changes result in Model (7) which presents symmetries similar to other clustering problems, any permutation of the loop assignments being equivalent. Constraint (7o) reduces the number of symmetric solutions by enforcing solutions with several actors decreasing with the loop index. Constraint (7k) is added to ensure that each actor cannot belong to two different loops.
@@ 3.5 Dantzig-Wolfe decomposition (MLCol)
Assumption 3.6 holds for Model (7).
@@
Furthermore, additional application-related constraints can directly be included in the loop, including, for instance, the fact that two coupled actors $(i,j)\in\mathcal{D}$ are both present or both absent from a loop:
@@
The extended formulation of Model (7) is essentially captured by the new constraints:
@@ 3.6 Benders decomposition of MLCol model (MLColExt)
s.t. $e_{ij}\leq D_{ij}\sum_{h\in\mathcal{H}}v_{h}a_{i}^{h}a_{j}^{h}$ $\forall i\in[n]\;\;$ $(\alpha_{ij})$ (11b)
$\sum_{j\in\mathcal{N}_{i}}e_{ij}+r_{i}-f_{i}-\sum_{j\in\mathcal{N}_{i}}e_{ji}=P_{i}-C_{i}$ $\forall i\in[n]\;\;$ $(\pi_{i})$ (11c)
$r_{i}\leq(1-\sum_{h\notin\mathcal{G}}v_{h}a_{i}^{h})Q$ $\forall i\in[n]\;\;$ $(\gamma_{i})$ (11d)
$\sum_{i\in[n]}r_{i}a_{i}^{h}\leq\sum_{i\in[n]}a_{i}^{h}v_{h}(P_{i}-C_{i})+Q(1-v_{h})$ $\forall i\in[n],h\in\mathcal{G}\;\;$ $(\nu_{h})$ (11e)
$\sum_{j\in\mathcal{N}_{i}}e_{ij}+r_{i}\leq P_{i}$ $\forall i\in[n]\;\;$ $(\tau_{i})$ (11f)
@@ 3.6.1 Dual and cuts
Dualisation of Problem (11) gives the following model, whose objective value provides cuts that can be applied to the main model to refine the chosen design of collective self-consumption loops: